OpenStudy (anonymous):

c=ax-bx Joseph claims that if a, b, and c are non-negative integers, then the equatoin has exactly one solution for x. Select all cases that show Joseph's claim is incorrect.

OpenStudy (dayakar):

c= ax-bx c= x(a-b) [ taking x as common] x(a-b) =c x= c/ (a-b) this is the solution and \[a-b \neq 0\]